OpenStudy (anonymous):

The solutions for a trig function (sin Thetha = 1/2) between [0, 2pi] are 7pi/6, and 11pi/6... why doesn't the answer include pi/6?

OpenStudy (primeralph):

maybe just a glitch

OpenStudy (anonymous):

I think he is right, did you find this in a book because it may be worth looking up to see if any mistakes were made in a specific edition. however, all multiples of pi/6 that fall in the first or second quadrant between 0 and 2pi should be included in the answer

OpenStudy (anonymous):

@jonsysibye @primeralph it was just what the professor wrote on the board

OpenStudy (primeralph):

well, from @jonsysibye there are other solutions.

OpenStudy (anonymous):

he may have made a mistake because all multiples of pi/6 in the first or second quadrant \[\sin \theta =1/2\]